OpenStudy (anonymous):

let y = 3x+ 10 - x^2, where x and y are positive whole no. find sum of all possible values of y.

OpenStudy (anonymous):

\[y =\left( x +2\right)\left( 5-x \right)\]

OpenStudy (amoodarya):



OpenStudy (anonymous):

y should be greater than or equal to zero. so x<5 given \[x \ge 0\] so substitute values between 0-5

OpenStudy (shubhamsrg):

Solve for x from the equation x^2 - 3x +(y-10) =0 => x = [ 3+- sqrt(49 - 4y) ]/2 First thing to take care of is 49-4y should be a perfect square for some whole number y -> 49 - 4y = m^2 (for some integer m) -> y = (49-m^2)/4 ->y =(7+m)(7-m)/4 clearly, m has to be odd number i.e. m^2 can only take values 1^2 ,3^2 ,5^2 .. keep in mind 49-4y >= 0 i.,e.y<= 12.25 max value of y can be 12. min can be anything for real number, but since y is a whole number, it can be minimum 0. for m^2 =1, y=12 m^2= 9 , y=10 m^2 = 25, y=6 m^2 = 49 , y=0 now check for each y, whether x is a whole number or not, x = [ 3+- sqrt(49 - 4y) ]/2 ofcorse it will be since sqrt(49 -4y) is always odd and 3 is also add, there addition or subtraction will yield an integer which is a multiple of 2. hence required ans = 12+10+6+0 = 38

OpenStudy (shubhamsrg):

28*

OpenStudy (anonymous):

thank u @shubhamsrg